Felicia is covered by a $180,000 group term life insurance policy and her daughter is the beneficiary


3. Felicia is covered by a $180,000 group term life insurance policy and her daughter is the beneficiary. Felicia's employer pays the entire cost of the policy for which the uniform annual premium is $8 per $1,000 of coverage. How much of this premium is taxable to Felicia?

 a. $0 b. $640 c. $1,040 d . $1,440 


4. Scott, age 22, is a full-time student at State College and a candidate for a bachelor's degree. During the current year, he received the following payments: State scholarship (tuition and books), $5,000: Loan from college financial aid office, $2,000; Cash support from parents, $9,000; Interest income from State of Texas bonds, $500; m father's employer upon father's death, $10,000. What is Scott's adjusted gross income for the year? 

a. $0 b. $5,000 c. $7,000 d. $10,000 e. $26,500
Answer #1

ANSWER: $1040

Explanation:

When employer provides life insurance to an employee , according to IRS it is considered as income in the hands of employee , which is subject to taxes. But this provision is applicable only if more than $50,000 in life insurance coverage is paid or payable by the employer . However there is an exemption for the premium cost for the first $50,000.

Annual premium = (8/1000)× 180000=$ 1,440

Exemption for 1st $50000=(8/1000)×50,000 =$400

Taxable life insurance premium=1440-400=$1040